OpenStudy (anonymous):

solve the following inequation: x+3≤(1-x^2)/(2-x)

OpenStudy (across):

"inequation" Haha! Well, why don't you try solving\[x+3=\frac{1-x^2}{2-x}\]instead?

OpenStudy (anonymous):

\[2<x\leq 5 \]

myininaya (myininaya):

\[x+3 \le \frac{1-x^2}{2-x}\] \[x+3-\frac{1-x^2}{2-x} \le 0\] \[\frac{(x+3)(2-x)-(1-x^2)}{2-x} \le 0\] \[\frac{(2x-x^2+6-3x)-(1-x^2)}{2-x} \le 0\] \[\frac{-x^2+x^2+2x-3x+6-1}{2-x} \le 0\] \[\frac{-x+5}{2-x} \le 0\] so the expression (-x+5)/(2-x) is zero when x=5 the expression (-x+5)/(2-x) is undefined when x=2 so --------|-----|---- 2 5 ------------------- -x+5| + | + | - | 2-x | + | - | - ============== + - + so the expression (-x+5)/(2-x) is less than zero when 2<x<5 ----------------------------------------------- final answer (-x+5)/(2-x) is less than or equal to when \[2 \le x <5\]

myininaya (myininaya):

oops

myininaya (myininaya):

\[2 < x \le 5\]

OpenStudy (anonymous):

Good work myininaya. I cheated and used the Mathematica "Reduce" function.

myininaya (myininaya):

lol

myininaya (myininaya):

thanks